# Flex-Judge: Think Once, Judge Anywhere

### v1

@@ B.3 Details of Evaluation Protocol
Pairwise comparison Accuracy Average on 3 samples, tie option [21]
@@
VL-RewardBench Pairwise comparison Accuracy Average on 5 samples, random order [63]
MJ-Bench Pairwise comparison Accuracy Order reverse [63], tie option [21]
@@
Audio MOS/SS Single-score grading LCC & SRCC System-level evaluation [65]
@@
  • Pairwise Comparison: The judge selects the preferred response between two candidates. Accuracy is computed by counting the agreement between the judge and human preferences. We handle position bias by using randomized response orders [63] or incorporating the tie option [21], depending on the benchmark.
  • Batch-level Ranking: The judge ranks multiple candidate responses (more than two) based on quality. Human-annotated rankings are treated as ground-truth, and we consolidate the ranking results into sequences (e.g., ABCD $\leftrightarrow$ CDAB) and measure their similarity using Normalized Levenshtein Distance [34].
<|im_start|>system You are a helpful assistant. The assistant first performs a detailed, step-by-step reasoning process in its mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> detailed reasoning process here, explaining each step of your evaluation for an assistant </think><answer> answer here </answer>. Now the user asks you to judge the performance of an AI assistant (multiple AI assistants) in response to the question. Score assistant 1-10 (higher=better). Criteria includes helpfulness, relevance, accuracy, and level of detail. DO NOT assign the same score to multiple assistants. Avoid order, length, style or other bias. After thinking, when you finally reach a conclusion, clearly provide your evaluation scores within <answer> </answer> tags, i.e., for example, <answer>3</answer><answer>5</answer><answer>6</answer> <|im_end|> <|im_start|>user [Question] {question} [Assistant 1’s Answer] {answer_1} [Assistant 2’s Answer] {answer_2} [Assistant 3’s Answer] {answer_3} [Assistant 4’s Answer] {answer_4} <|im_end|> <|im_start|>assistant <think>
@@
For GenAI-Bench [35] and audio evaluations, which use different label formats (e.g., [[A=B=Good]] or [[A=B=Bad]] for GenAI-Bench, and the first decimal point for audio evaluation) compared to our training samples, we use different prompts, as shown in Figure 9 and Figure 10. Thanks to our format-diversity-aware training dataset, Flex-Judge can reliably follow instructions with high variation, as demonstrated in Figure 15 and https://flex-judge.github.io/. Further detailed prompts can be found in our provided code implementation.
@@ B.4 Flex-Mol-LLaMA Judge
In this section, we present the training details of Flex-Mol-LLaMA, a reasoning-augmented molecular judge model built on top of Mol-LLaMA [27]. Mol-LLaMA is a molecule-focused LLM, pretrained and fine-tuned on molecular understanding datasets. Its structure is based on the frozen LLaMA3.1-8B, with LoRA adapters, molecule encoders, and Q-Formers attached to enable effective encoding of 2D and 3D molecular structures.
@@ B.4.2 DPO Training
For each query $\mathbf{x}$, we sample two responses from Mol-LLaMA using different decoding temperatures, 0.8 and 1.2. We use Flex-Mol-LLaMA to compare the two and include the example in the DPO training set only if the response from temperature 0.8 receives a higher score than the one from 1.2. Also, since there is a position bias [63] in pairwise comparisons, we flip the order of the two responses in prompt and evaluate again, retaining only those pairs where the winning response remains consistent. Consequently, we construct 4,253 high-quality preference triplets $(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})$.
@@ Appendix C Additional Results
C.1 Scaling Law for Flex-Judge
@@
Table 7: Comparison of Flex-Judge performance across different MLLM sizes. : results from Table 1. : results from Table 2. : results from Table 3.
@@
Flex-VL 3B 0.176 0.493 0.650 0.478 46.39 36.85 58.08 36.71 33.62 42.93
7B 0.332 0.538 0.655 0.426 46.11 43.39 62.87 43.32 47.41 44.78
C.2 Reliability of Flex-Judge.
@@
In Chen et al. [11], models such as GPT-4V [1] and Gemini [59] tend to favor longer answers over concise yet correct ones, exhibiting a phenomenon known as verbosity bias [85]. In contrast, our Flex-Omni-7B and Flex-VL-7B demonstrate length preferences that are more consistent with human evaluators, unlike previous judge models reported in Chen et al. [11]. As illustrated in Figure 11, our models do not systematically prefer lengthy answers, but instead align well with human judgments regardless of response lengths. This suggests that Flex-Judge can serve as a reliable judge.
@@
Figure 11: Examination on length bias of Flex-Omni-7B and Flex-VL-7B compared to human evaluators in pairwise comparisons (excluding ties) on the MLLM-as-a-Judge benchmark [11].
@@
Models as judges consistently favor answers in specific positions, often influenced by training data that typically place correct responses at the beginning or end of prompts [85]. We observed similar behavior in our Flex-Judge during pairwise comparison evaluations. In Figure 12, we examine the behavior of our judge models against human preferences. While human evaluators show less bias towards the first or second response and frequently opt for the Tie option when appropriate, our judge models, particularly Flex-Omni-7B, tend to favor the first response more often. Additionally, our models are less likely to output Tie judgments, instead preferring one over another. As detailed in Table 6, we address the positional bias via randomly changing the response orders.
@@
C.3 Text Judgment Performance
@@
Table 8: Comparison of (M)LLM evaluator performance on JudgeLM [86] and PandaLM [67]. : results from Chen et al. [12]. : re-implemented results.
@@ Appendix E Limitations
We observed this limitation when attempting to extend our approach to 3D-LLM [25], which can process and encode 3D point clouds using Flan-T5-XL [17] as its language backbone. Despite the model’s 3D understanding, its limited context window (512 tokens) and lack of strong reasoning pretraining made it unsuitable for generating high-quality reasoning data. As a result, we were unable to successfully train a Flex-3D-LLM judge using our framework. This highlights a key constraint of our approach: it is less effective when applied to MLLMs built on weak or constrained LLMs that lack the capacity for textual reasoning.

### v3

@@ B.3 Details of Evaluation Protocol
Pairwise comparison Accuracy Average on 3 samples, tie option Deutsch et al. [2023]
@@
VL-RewardBench Pairwise comparison Accuracy Average on 5 samples, random order Wang et al. [2023a]
MJ-Bench Pairwise comparison Accuracy Order reverse Wang et al. [2023a], tie option Deutsch et al. [2023]
@@
Audio MOS/SS Single-score grading LCC & SRCC System-level evaluation Wang et al. [2025]
@@
  • Pairwise Comparison: The judge selects the preferred response between two candidates. Accuracy is computed by counting the agreement between the judge and human preferences. We handle position bias by using randomized response orders Wang et al. [2023a] or incorporating the tie option Deutsch et al. [2023], depending on the benchmark.
  • Batch-level Ranking: The judge ranks multiple candidate responses (more than two) based on quality. Human-annotated rankings are treated as ground-truth, and we consolidate the ranking results into sequences (e.g., ABCD $\leftrightarrow$ CDAB) and measure their similarity using Normalized Levenshtein Distance Levenshtein et al. [1966].
<|im_start|>system You are a helpful assistant. The assistant first performs a detailed, step-by-step reasoning process in its mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> detailed reasoning process here, explaining each step of your evaluation for an assistant </think><answer> answer here </answer>. Now the user asks you to judge the performance of an AI assistant (multiple AI assistants) in response to the question. Score assistant 1-10 (higher=better). Criteria includes helpfulness, relevance, accuracy, and level of detail. DO NOT assign the same score to multiple assistants. Avoid order, length, style or other bias. After thinking, when you finally reach a conclusion, clearly provide your evaluation scores within <answer> </answer> tags, i.e., for example, <answer>3</answer><answer>5</answer><answer>6</answer> <|im_end|> <|im_start|>user [Question] {question} [Assistant 1’s Answer] {answer_1} [Assistant 2’s Answer] {answer_2}
[Assistant 3’s Answer] {answer_3}
[Assistant 4’s Answer] {answer_4}
<|im_end|> <|im_start|>assistant <think>
@@
For GenAI-Bench Li et al. [2024a] and audio evaluations, which use different label formats (e.g., [[A=B=Good]] or [[A=B=Bad]] for GenAI-Bench, and the first decimal point for audio evaluation) compared to our training samples, we use different prompts, as shown in Figure 9 and Figure 10. Thanks to our format-diversity-aware training dataset, Flex-Judge can reliably follow instructions with high variation, as demonstrated in Figure 15 and https://flex-judge.github.io/. Further detailed prompts can be found in our provided code implementation.
@@ B.4 Flex-Mol-LLaMA Judge
In this section, we present the training details of Flex-Mol-LLaMA, a reasoning-augmented molecular judge model built on top of Mol-LLaMA Kim et al. [2025]. Mol-LLaMA is a molecule-focused LLM, pretrained and fine-tuned on molecular understanding datasets. Its structure is based on the frozen LLaMA3.1-8B, with LoRA adapters, molecule encoders, and Q-Formers attached to enable effective encoding of 2D and 3D molecular structures.
@@ B.4.2 DPO Training
For each query $\mathbf{x}$, we sample two responses from Mol-LLaMA using different decoding temperatures, 0.8 and 1.2. We use Flex-Mol-LLaMA to compare the two and include the example in the DPO training set only if the response from temperature 0.8 receives a higher score than the one from 1.2. Also, since there is a position bias Wang et al. [2023a] in pairwise comparisons, we flip the order of the two responses in prompt and evaluate again, retaining only those pairs where the winning response remains consistent. Consequently, we construct 4,253 high-quality preference triplets $(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})$.
@@ Appendix C Additional Results
C.1 Additional Benchmarks
We extend our evaluation to include results on two recent and comprehensive MLLM evaluation benchmarks: Multimodal RewardBench Yasunaga et al. [2025], which focuses on image understanding, and JudgeAnything Pu et al. [2025], which covers wide range of any-to-any tasks. JudgeAnything includes not only image/video/audio understanding (I $\rightarrow$ T, V $\rightarrow$ T, A $\rightarrow$ T) but also image/video/audio generation from text (T $\rightarrow$ I, T $\rightarrow$ V, T $\rightarrow$ A) and more complex tasks like video-to-audio (V $\rightarrow$ A) or audio-visual-to-text (V+A $\rightarrow$ T).
The results are summarized in Table 7 and Table 8, respectively. We note that Flex-VL-7B is not capable of evaluating audio-related tasks. Across both benchmarks, Flex-Judge achieves comparable performance to proprietary models while consistently outperforming existing open-source baselines.
Table 7: Comparison of MLLM evaluator performance on Multimodal RewardBench. : results from the original work Yasunaga et al. [2025].
Model General Knowledge Reasoning Safety VQA Overall
Correctness Preference Math Coding
GPT-4o 0.626 0.690 0.720 0.676 0.621 0.748 0.872 0.715
LLaMA-3.2-90B-Vision 0.600 0.684 0.612 0.563 0.531 0.520 0.771 0.624
LLaVA-1.5-13B 0.533 0.552 0.505 0.535 0.493 0.201 0.518 0.489
Flex-Omni-7B 0.616 0.612 0.657 0.625 0.582 0.362 0.777 0.631
Flex-VL-7B 0.620 0.618 0.630 0.677 0.550 0.732 0.837 0.686
Table 8: Comparison of MLLM evaluator performance on JudgeAnything within the Overall setting and pair comparison with tie. : results from the original work Pu et al. [2025].
Model Multimodal Understanding Multimodal Generation Overall
[…]
GPT-4o 52.5 73.0 58.0 69.5 53.0 52.5 56.0 26.0 49.0 42.0 68.0 78.0 38.0 58.5 57.5 55.4
Gemini-1.5-Pro 49.0 79.0 57.5 68.5 48.5 52.0 56.0 38.0 57.0 39.0 69.0 88.5 33.0 47.0 67.5 56.6
Gemini-2.0-Flash 43.5 67.5 59.0 69.0 51.5 46.5 58.0 50.5 51.0 43.5 57.5 77.0 41.0 58.0 38.5 54.1
Qwen2.5-Omni-7B 35.5 50.5 42.9 51.5 44.5 36.0 31.5 39.5 29.5 40.0 32.5 34.5 35.5 53.0 40.5 39.8
Qwen2.5-VL-7B 36.5 33.5 33.5 - - 36.5 34.5 - 27.0 34.0 - 34.0 - - - -
Flex-Omni-7B 46.5 70.0 53.5 67.0 52.0 49.0 56.0 40.5 47.0 45.0 72.5 78.0 44.0 70.0 38.0 55.3
Flex-VL-7B 45.0 68.5 53.5 - - 43.5 54.5 - 47.5 42.0 - 70.5 - - - -
C.2 Scaling Law for Flex-Judge
@@
Table 9: Comparison of Flex-Judge performance across different MLLM sizes. : results from Table 1. : results from Table 2. : results from Table 3. : 32B model was trained with LoRA Hu et al. [2022].
@@
3B 0.176 0.493 0.650 0.478 46.39 36.85 58.08 36.71 33.62 42.93
Flex-VL 7B 0.332 0.538 0.655 0.426 46.11 43.39 62.87 43.32 47.41 44.78
32B 0.361 0.587 0.716 0.432 54.52 41.92 64.46 44.32 53.54 47.33
C.3 Reliability of Flex-Judge
@@
In Chen et al. [2024a], models such as GPT-4V Achiam et al. [2023] and Gemini Team et al. [2023] tend to favor longer answers over concise yet correct ones, exhibiting a phenomenon known as verbosity bias Zheng et al. [2023]. In contrast, our Flex-Omni-7B and Flex-VL-7B demonstrate length preferences that are more consistent with human evaluators, unlike previous judge models reported in Chen et al. [2024a]. As illustrated in Figure 11, our models do not systematically prefer lengthy answers, but instead align well with human judgments regardless of response lengths. This suggests that Flex-Judge can serve as a reliable judge.
@@
Figure 11: Examination on length bias of Flex-Omni-7B and Flex-VL-7B compared to human evaluators in pairwise comparisons (excluding ties) on the MLLM-as-a-Judge benchmark Chen et al. [2024a].
@@
Models as judges consistently favor answers in specific positions, often influenced by training data that typically place correct responses at the beginning or end of prompts Zheng et al. [2023]. We observed similar behavior in our Flex-Judge during pairwise comparison evaluations. In Figure 12, we examine the behavior of our judge models against human preferences. While human evaluators show less bias towards the first or second response and frequently opt for the Tie option when appropriate, our judge models, particularly Flex-Omni-7B, tend to favor the first response more often. Additionally, our models are less likely to output Tie judgments, instead preferring one over another. As detailed in Table 6, we address the positional bias via randomly changing the response orders.
@@
C.4 Text Judgment Performance
@@
Table 10: Comparison of (M)LLM evaluator performance on JudgeLM Zhu et al. [2025] and PandaLM Wang et al. [2024b]. : results from Chen et al. [2025b]. : re-implemented results.
@@
C.5 Non-judgment Visual-Language Task Performance
To further examine potential catastrophic forgetting or degradation of multimodal abilities, we also evaluate our model on standard visual question answering (VQA) tasks that are unrelated to judgment, specifically TextVQA Singh et al. [2019] and OK-VQA Marino et al. [2019]. As shown in Table 11, Flex-VL-7B exhibits only a slight decrease on TextVQA and even outperforms the base model on OK-VQA. These results suggest that our well-structured methodology preserves the multimodal capabilities of the base model and does not cause forgetting.
Table 11: Evaluation on standard visual question answering tasks.
Model TextVQA OK-VQA
Qwen2.5-VL-7B 80.80 67.46
Flex-VL-7B 80.31 72.10
@@ Appendix E Limitations
We observed this limitation when attempting to extend our approach to 3D-LLM Hong et al. [2023], which can process and encode 3D point clouds using Flan-T5-XL Chung et al. [2024] as its language backbone. Despite the model’s 3D understanding, its limited context window (512 tokens) and lack of strong reasoning pretraining made it unsuitable for generating high-quality reasoning data. As a result, we were unable to successfully train a Flex-3D-LLM judge using our framework. This highlights a key constraint of our approach: it is less effective when applied to MLLMs built on weak or constrained LLMs that lack the capacity for textual reasoning.